### v1

@@ Abstract
[…]
@@ 1 Introduction
Convolutional Neural Networks (CNNs) are a cornerstone of computer vision algorithms, largely owing to their proficiency in managing spatial relationships and maintaining invariance to input translations. Their widespread success in object recognition tasks can be attributed to advantageous inductive biases, such as translation equivalence, which enable them to effectively identify and process visual patterns. The use of convolutions in neural networks for vision-related tasks, a methodology initiated with the early implementations [1, 2], and was greatly popularized by the work of Krizhevsky et al. [3], has led to the scenario where nearly most vision networks rely on convolutional layers [4, 5, 6, 7, 8]. These convolutional layers have been further improved with innovations such as residual connections [5]. Extensive use has led to detailed empirical [9] and analytical evaluations of convolutional networks [10, 11].
@@
Recent advancements, however, have shown the potential of transformers regarding to their self-attention mechanisms that find the global features of the data which provide a more holistic view of the data, reduce inductive bias, and exhibit a high degree of scalability and flexibility. These factors collectively enhance the model’s ability to generalize better during testing. After the tremendous success in Natural Language Processing (NLP) tasks [12, 13, 14], transformers are now being actively integrated into computer vision tasks [15, 16, 17, 18, 19]. Attention-based architectures known as transformers, gaining traction in both pure vision and vision-language tasks [20, 21, 22, 23, 24, 25]. Vision Transformers (ViT) [16], stands out as a key development in this area, applying a self-attention-based mechanism to sequences of image patches. It achieves competitive performance on the challenging ImageNet classification task [26], compared to CNNs. Researchers discovered that existing CNN architectures exhibit limited generalization capabilities when confronted with distribution shift scenarios [27, 28]. Subsequent research, as seen in works like [29, 30], has further expanded the capabilities of transformers, demonstrating impressive performance across various visual benchmarks. This includes applications in different benchmarks including COCO (Common Objects in Context) dataset in object detection and instance segmentation [31], as well as ADE20K dataset for semantic segmentation [32].
@@
As ViTs gain popularity, it becomes crucial to examine the characteristics of the representations they learn [33]. This is important in areas like autonomous driving [34, 35], robotics [36], and healthcare [37], where the trustworthy and reliability of these systems are crucial. Recent studies delve into evaluating ViTs’ robustness, focusing not just on standard metrics like accuracy and computational cost, but also on their intrinsic impact on model robustness and generalization, especially in handling Distribution Shifts. In conventional training and testing scenarios, it is assumed that data are independent and identically distributed (IID). However, this assumption often doesn’t reflect real-world scenarios. Therefore, exploring the potential of ViTs as the modern vision networks, to adapt to target domains, and generalize and perform well on unseen data, becomes a crucial aspect of machine learning models [38].
@@
Building on this foundation, research efforts have been directed at enhancing ViTs’ generalization capabilities through various methodologies, delving into both DA and DG strategies, and integrating ViTs into the broader deep learning framework. Comparative analyses of ViTs and high-performing CNNs reveal distinct advantages attributable to ViTs. A key differentiation is the dynamic nature of weight computation in ViTs through the self-attention mechanism, contrasting with the static weights learned by CNNs during training. This attribute provides ViTs with a more flexible and adaptable ability [46]. ViTs employ multi-head self-attention to intricately parse and interpret contextual information within images, thereby excelling in scenarios involving occlusions, domain variations, and perturbations. They demonstrate remarkable robustness, effectively maintaining accuracy despite image modifications. Furthermore, ViTs exhibit a reduced texture bias compared to CNNs, favoring shape recognition which aligns more closely with human visual processing. This proficiency in discerning overall shapes facilitates accurate image categorization without relying on detailed, pixel-level analysis. ViTs’ ability to merge various features for image classification enhances their performance across diverse datasets, proving advantageous in both conventional and few-shot learning settings where the model is trained with only a few examples [33, 47, 48]. Refer to Figure 1 for the challenges prevalent in images. These challenges include severe occlusions, adversarial perturbations, patch permutations, and domain shifts effectively addressed through the flexibility and dynamic receptive fields of self-attention mechanisms [33]. Unlike CNNs, which primarily focus on texture [49], ViTs concentrate more on object shape, enhancing their resistance to texture shifts and benefiting shape recognition, and they are adept at propagating spatial information, an advantage for tasks such as detection and segmentation [33, 48].
@@
In our comprehensive review, the first of its kind to integrate ViTs for Domain Adaptation (DA) and Domain Generalization (DG) scenarios, we have meticulously examined how ViTs adapt to distribution shifts. This study delves into the fundamentals, architecture, and key components of ViTs, offering a unique categorization and analysis of their role in domain adaptation and generalization. Within the context of DA, we cover feature-level, instance-level, model-level, and hybrid approaches, alongside a variety of strategies to enhance domain adaptation through ViTs. For DG, our exploration includes multi-domain learning, meta-learning approaches, regularization techniques, data augmentation strategies, and further methods to boost domain generalization. A significant portion of our review is dedicated to showcasing the applications of ViTs beyond image recognition, such as semantic segmentation, action recognition, face analysis, medical imaging, and other emerging fields. This broad spectrum of applications highlights the versatility and potential of ViTs across the vast landscape of computer vision. In our discussion section, we delve into the initial development challenges associated with ViTs, aiming to equip researchers with insightful findings that could steer the direction of future investigations in this area. Furthermore, we outline prospective research paths.
As the field of computer vision advances, especially with the introduction of ViTs, this survey emerges as a crucial tool for researchers and practitioners alike. We aspire that our findings will stimulate further investigation and innovation in leveraging ViTs for Domain Adaptation and Generalization, thereby overcoming current hurdles and paving new paths in this ever-evolving research domain.
@@ 2.4 Advantages of Vision Transformers compared to CNNs backbones
Exploiting Long-Range Dependencies, the Power of Attention Mechanisms in ViTs: The attention mechanism within ViTs effectively captures long-range dependencies in the input data. This modeling of inter-token relationships facilitates a more comprehensive global context, a significant step beyond the local processing capabilities of CNNs [78, 17, 79]. Additionally, the attention mechanism provides insight into the focus areas of the model during input processing, acting as a built-in saliency map [80].
@@ 3 Vision Transformers in Domain Adaptation and Domain Generalization
Figure 4: Adapting ViTs to distribution shifts involves categorizing the studies based on methods and strategies developed to enhance ViTs’ robustness and generalization in the face of varying distributions.
@@ 3.1.1 Feature-Level Adaptation
Feature-Level adaptation encompasses techniques where ViTs are adapted at the feature extraction level. It often involves aligning feature distributions between source and target domains.
Researchers proposed a Domain-Oriented Transformer (DOT) to address the challenges faced by conventional UDA techniques for domain discrepancies. Traditional methods often encounter difficulties when attempting to align domains, which can compromise the discriminability of the target domain when classifiers are biased towards the source data. To overcome these limitations, DOT employs feature alignment across two distinct spaces, each specifically designed for one of the domains. It leverages separate classification tokens and classifiers for each domain. This approach ensures the preservation of domain-specific discriminability while effectively capturing both domain-invariant and domain-specific information. It achieves this through a combination of contrastive-based alignment and source-guided pseudo-label. The novel DOT method is introduced in [87].
@@ 3.1.2 Instance-Level Adaptation
In this section, we shift our focus to techniques that adjust the importance of individual data points or modify how they are selected during training. These methods give priority to specific instances that more closely reflect the characteristics of the target domain, an approach known as instance-level adaptation. This involves selecting or weighing certain data points more heavily during the training process to ensure the model learns features that are relevant to the new domain.
One notable instance of instance-level adaptation is addressed in the study by [96], which deals with the challenge of Source Free Open Set Domain Adaptation (SF-OSDA). This scenario involves adapting a pre-trained model, initially trained on an inaccessible source dataset, to an unlabeled target dataset that includes open set samples, or data points that do not belong to any class seen during training. The primary technique involves leveraging a self-supervised ViT, which learns directly from the target domain to distill knowledge. A crucial element of their method is a unique style-based data augmentation technique, designed to enhance the training of the ViT within the target domain by providing a richer, more contextually diverse set of training data. This leads to the creation of embeddings with rich contextual information.
@@ 3.1.3 Model-Level Adaptation
Model-level adaptation includes the development of specialized ViT architectures or layers, designed to enhance adaptability to domain shifts. Key examples include attention mechanisms that focus on domain-invariant features. This segment of the analysis highlights various strategies employed in the adaptation of the entire model architecture of ViTs, signifying a key aspect of the study’s in-depth review.
@@ 3.1.4 Hybrid Approaches
Hybrid approaches combine features, instance, and model-based methods for more robust adaptation. CDTRANS [84] introduces a hybrid domain adaptation approach via a triple-branch transformer that combines feature-level and model-level adaptations. It incorporates a novel cross-attention module along with self-attention mechanisms within its architecture. The design includes separate branches for source and target data processing and a third for aligning features from both domains. This setup enables simultaneous learning of domain-specific and domain-invariant representations, showing resilience against label noise. Additionally, the paper proposes a two-way center-aware labeling algorithm for the target domain, utilizing a cross-domain similarity matrix to enhance pseudo-label accuracy and mitigate noise impact. This approach effectively merges feature acquisition with alignment, showcasing a sophisticated method for handling domain adaptation challenges. TVT (Transferable Vision Transformer) [85] explores the use of ViT in UDA. It examines ViT’s transferability compared to CNNs and proposes the TVT framework, which includes a Transferability Adaptation Module (TAM) and a Discriminative Clustering Module (DCM). TVT integrates various domain adaptation approaches. It employs model-level adaptations by leveraging ViTs, and optimizing them for UDA. Simultaneously, it involves feature-level adaptation through the TAM, which enhances feature representations for better alignment between source and target domains. Furthermore, instance-level strategies are incorporated via the Discriminative Clustering Module (DCM), focusing on the diversification and discrimination of patch-level features. This multifaceted approach, combining model, feature, and instance-level adaptations, exemplifies a hybrid strategy in domain adaptation.
@@
Category
Method
Design Highlights
Loss Functions
Publication
Feature-level Adaptation
DOT [87]
Domain-Oriented Transformer with dual classifiers utilizes individual tokens for source/target domain adaptation and domain-specific learning, enhanced by source-guided pseudo-label refinement
Cross-entropy, Contrastive, Representation Difference
ICMR 2022
SUDA [92]
Enhancing visual task generalization, leveraging Spectrum Transformer for domain alignment and multi-view learning to optimize mutual information
Supervised, Discrepancy, Unsupervised Similarity
CVPR 2022
Model-level Adaptation
GE-ViT [38]
Leveraging ViTs’ inductive biases towards shapes and structures, combined with adversarial and self-supervised learning techniques
Cross-Entropy, Entropy, Adversarial
CVPR 2022
TFC [99]
Integrates CNN and ViTs for both capturing of local and global features, achieving distinct foreground-background separation, further refined by an uncertainty-based accuracy enhancement
Cross-Entropy, Adversarial, Uncertainty Penalization
IEEE-TCSS 2022
Hybrid Approaches
CDTRANS [84]
Triple-branch Transformer with shared weights aligns source/target data, using self/cross-attention and center-aware labeling for enhanced pseudo-label accuracy and noise distillation
Cross-entropy, Distillation
ICLR 2022
TVT [85]
Enhances feature learning via adversarial adaptation, using transferability and clustering modules for robust knowledge transfer, and fine-tunes attention for precise domain discrimination
Cross-Entropy, Adversarial, Patch-Level Discriminator
WACV2023
Hybrid Approaches
SSRT [86]
Utilizes adversarial adaptation and enhances model accuracy by refining through perturbed data feedback, incorporating a robust training approach to minimize KL divergence
Cross-Entropy, Self-Refinement, Domain Adversarial
CVPR2022
PMTrans [88]
Employs a module based on transformer architecture for domain representation, leveraging game theory for patch sampling, and combines source/target data to optimize label accuracy with attention-driven adjustments
Cross-Entropy, Semi-Supervised Mixup
CVPR 2023
SMTDA [102]
Enhances domain adaptation by maximizing nuclear-norm consistency and distilling knowledge via MixUp, refining pseudo labels for broadened generalization across multiple targets
Cross-Entropy, Nuclear-Norm, Knowledge Distillation
WACV 2023
UniAM [101]
Utilizes universal and compressive attention for precise alignment and class separation, achieving categorization without prior label knowledge
Cross-Entropy, Adversarial, Source-Target Contrastive
ICCV 2023

### v2

@@ Abstract
Vision Transformers, Domain Adaptation, Domain Generalization, Distribution Shifts
@@ 1 Introduction
Convolutional Neural Networks (CNNs) are a cornerstone of computer vision algorithms, largely owing to their proficiency in managing spatial relationships and maintaining invariance to input translations. Their widespread success in object recognition tasks can be attributed to advantageous inductive biases, such as translation equivalence, which enable them to effectively identify and process visual patterns. The foundational concept of using convolutions in neural networks was initiated by Fukushima’s development of the Neocognitron [1], a model that introduced the idea of a shift-invariant architecture. This idea was further advanced by LeCun et al., who applied gradient-based learning to document recognition, showcasing the practical applicability of CNNs [2]. The significant breakthrough in CNNs came with Krizhevsky et al., whose work on ImageNet classification popularized deep convolutional networks [3]. Following this, developments such as Szegedy et al.’s deeper convolutional networks [4], He et al.’s introduction of residual networks [5], and Huang et al.’s densely connected networks [6] have each contributed unique architectural improvements that enhance model robustness and accuracy. Recent studies like those by Hsieh et al. [7] and Tan and Le [8] continue to explore the limits of CNN efficiency and robustness, further solidifying the central role of convolutional layers in modern vision networks. These convolutional layers have been further improved with innovations such as residual connections [5]. Extensive use has led to detailed empirical [9] and analytical evaluations of convolutional networks [10, 11].
Recent advancements, however, have shown the potential of transformers regarding to their self-attention mechanisms that find the global features of the data which provide a more holistic view of the data [12], reduce inductive bias, and exhibit a high degree of scalability and flexibility. These factors collectively enhance the model’s ability to generalize better during testing. After their tremendous success in Natural Language Processing (NLP) tasks, transformers are now being actively integrated into computer vision tasks. Pioneering works like Vaswani et al. [13] introduced transformers, showcasing their efficiency in handling long-range dependencies in data. This approach was extended to language understanding by Devlin et al. [14] with BERT, which dramatically improved the performance of NLP tasks. Brown et al. [15] further demonstrated the capability of transformers in NLP with their few-shot learning approaches. In the realm of computer vision, Chen et al. [16] and Dosovitskiy et al. [17] adapted transformer architectures to manage spatial hierarchies in images, leading to significant advancements in image segmentation and recognition tasks. Touvron et al. [18] explored training data-efficient image transformers, which optimize the transformer architecture for better performance with limited data. Khan et al. [19] provided a comprehensive survey on the application of transformers in vision, encapsulating various models and methodologies that have evolved over time. Adaptformer by Chen et al. [20] adapts ViTs for scalable visual recognition, enhancing their adaptability and efficiency across different scales. In terms of integrating vision and language tasks, works like VideoBERT by Sun et al. [21] and ViLBERT by Lu et al. [22] have been foundational, developing joint models that learn correlated features between video and text. LXMERT by Tan et al. [23] and UNITER by Chen et al. [24] further refine these approaches, improving the cross-modal understanding necessary for complex tasks involving both vision and language. Finally, Radford et al. [25] explored the use of transformers to develop visual models that can be trained using only natural language descriptions, rather than traditional image labels. This approach leverages the rich contextual information available in language to enhance the model’s ability to understand and generalize across different visual and textual modalities. By doing so, they are advancing the capacity of models to perform tasks in more diverse and complex environments, effectively bridging the gap between vision and language.
@@
Vision Transformers (ViT) [17], stands out as a key development in this area, applying a self-attention-based mechanism to sequences of image patches. It achieves competitive performance on the challenging ImageNet classification task [26], compared to CNNs. Researchers discovered that existing CNN architectures exhibit limited generalization capabilities when confronted with distribution shift scenarios [27, 28]. Subsequent research, as seen in works like [29, 30], has further expanded the capabilities of transformers, demonstrating impressive performance across various visual benchmarks. This includes applications in different benchmarks including COCO (Common Objects in Context) dataset in object detection and instance segmentation [31], as well as ADE20K dataset for semantic segmentation [32].
@@
As ViTs gain popularity, it becomes crucial to examine the characteristics of the representations they learn [33]. This is important in areas like autonomous driving [34, 35], robotics [36], and healthcare [37, 38], where the trustworthy and reliability of these systems are crucial. Recent studies delve into evaluating ViTs’ robustness, focusing not just on standard metrics like accuracy and computational cost, but also on their intrinsic impact on model robustness and generalization, especially in handling Distribution Shifts. In conventional training and testing scenarios, it is assumed that data are independent and identically distributed (IID). However, this assumption often doesn’t reflect real-world scenarios. Therefore, exploring the potential of ViTs as the modern vision networks, to adapt to target domains, and generalize and perform well on unseen data, becomes a crucial aspect of machine learning models [39].
@@
Building on this foundation, research efforts have been directed at enhancing ViTs’ generalization capabilities through various methodologies, delving into both DA and DG strategies, and integrating ViTs into the broader deep learning framework. Comparative analyses of ViTs and high-performing CNNs reveal distinct advantages attributable to ViTs. A key differentiation is the dynamic nature of weight computation in ViTs through the self-attention mechanism, contrasting with the static weights learned by CNNs during training. This attribute provides ViTs with a more flexible and adaptable ability [47]. ViTs employ multi-head self-attention to intricately parse and interpret contextual information within images, thereby excelling in scenarios involving occlusions, domain variations, and perturbations. They demonstrate remarkable robustness, effectively maintaining accuracy despite image modifications [48]. Furthermore, ViTs exhibit a reduced texture bias compared to CNNs, favoring shape recognition which aligns more closely with human visual processing. This proficiency in discerning overall shapes facilitates accurate image categorization without relying on detailed, pixel-level analysis. ViTs’ ability to merge various features for image classification enhances their performance across diverse datasets, proving advantageous in both conventional and few-shot learning settings where the model is trained with only a few examples [33, 49, 50]. Refer to Figure 1 for the challenges prevalent in images. These challenges include severe occlusions, adversarial perturbations, patch permutations, and domain shifts effectively addressed through the flexibility and dynamic receptive fields of self-attention mechanisms [33]. Unlike CNNs, which primarily focus on texture [51], ViTs concentrate more on object shape, enhancing their resistance to texture shifts and benefiting shape recognition, and they are adept at propagating spatial information, an advantage for tasks such as detection and segmentation [33, 50].
@@
In our comprehensive review, the first of its kind to explore the potential of ViTs in DA and DG scenarios, we have meticulously examined how ViTs adapt to distribution shifts. This study delves into the fundamentals, architecture, and key components of ViTs, offering a unique categorization and analysis of their role in both the theoretical aspects and practical implementations of DA and DG. In this review paper, we explored all the existing papers in this field and developed our own categorizations for the research. Within the context of DA, we categorized the research into feature-level, instance-level, model-level, and hybrid approaches. For DG, our categorization includes multi-domain learning, meta-learning approaches, regularization techniques, and data augmentation strategies. After the first categorization, we found that there are many studies using ViTs in different categories for DA and DG. The research is somewhat sparse, with researchers applying these modern vision networks in various DA and DG strategies, making it challenging to categorize this diverse research. To address this, we introduced another categorization, providing tables including each study that use different methods. We defined the methods they are using and then, for each study, specified these methods. This dual categorization approach is particularly useful because most of the research employs hybrid methods. A significant portion of our review is dedicated to showcasing the applications of ViTs beyond image recognition, such as semantic segmentation, action recognition, face analysis, medical imaging, and other emerging fields. This broad spectrum of applications highlights the versatility and potential of ViTs across the vast landscape of computer vision. In our discussion section, we delve into the initial development challenges associated with ViTs, aiming to equip researchers with insightful findings that could steer the direction of future investigations in this area. Furthermore, we outline prospective research paths.
As the field of computer vision advances, especially with the introduction of ViTs, this survey emerges as a promising source for researchers and practitioners alike. We aspire that our findings will stimulate further investigation and innovation in leveraging ViTs for Domain Adaptation and Generalization, thereby overcoming current hurdles and paving new paths in this ever-evolving research domain.
@@ 2.4 Advantages of Vision Transformers compared to CNNs backbones
Exploiting Long-Range Dependencies, the Power of Attention Mechanisms in ViTs: The attention mechanism within ViTs effectively captures long-range dependencies in the input data. This modeling of inter-token relationships facilitates a more comprehensive global context, representing a significant advancement beyond the local processing capabilities of CNNs [80, 18, 81] and more recent works [82]. Additionally, the attention mechanism provides insight into the focus areas of the model during input processing, acting as a built-in saliency map [83].
@@ 3 Vision Transformers in Domain Adaptation and Domain Generalization
Figure 4: Our categorization of studies on adapting vision transformers to handle distribution shifts in domain adaptation and domain generalization approaches.
@@ 3.1.1 Feature-Level Adaptation
Feature-level adaptation involves aligning the feature distributions between the source and target domains to ensure that the features learned by the model in the source domain are applicable to the target domain. This approach is particularly effective in addressing the domain shift problem by transforming the feature space of the source domain to closely match that of the target domain. Techniques such as Domain-Oriented Transformer (DOT), TRANS-DA, and Spectral UDA (SUDA) have shown promising results by employing various strategies like adversarial training, feature matching, and domain-specific normalization. By aligning feature distributions, these methods help to reduce the discrepancy between domains, thereby improving the model’s performance on the target domain.
Researchers proposed a Domain-Oriented Transformer to address the challenges faced by conventional UDA techniques for domain discrepancies. Traditional methods often encounter difficulties when attempting to align domains, which can compromise the discriminability of the target domain when classifiers are biased towards the source data. To overcome these limitations, DOT employs feature alignment across two distinct spaces, each specifically designed for one of the domains. It leverages separate classification tokens and classifiers for each domain. This approach ensures the preservation of domain-specific discriminability while effectively capturing both domain-invariant and domain-specific information. It achieves this through a combination of contrastive-based alignment and source-guided pseudo-label. The novel DOT method is introduced in [90].
@@ 3.1.2 Instance-Level Adaptation
Instance-level adaptation involves selecting or weighting specific data points (instances) more heavily during training to ensure that the model learns features relevant to the target domain. This approach can significantly improve the model’s performance on the target domain by prioritizing instances that reflect the characteristics of the target data. Techniques such as Source Free Open Set Domain Adaptation (SF-OSDA), style-based data augmentation, and clustering are commonly used in instance-level adaptation to refine and enhance the training process. By focusing on relevant instances, these methods help to reduce the impact of domain shift and improve the model’s generalization capabilities.
One notable instance of instance-level adaptation is addressed in the study by [99], which deals with the challenge of Source Free Open Set Domain Adaptation. This scenario involves adapting a pre-trained model, initially trained on an inaccessible source dataset, to an unlabeled target dataset that includes open set samples, or data points that do not belong to any class seen during training. The primary technique involves leveraging a self-supervised ViT, which learns directly from the target domain to distill knowledge. A crucial element of their method is a unique style-based data augmentation technique, designed to enhance the training of the ViT within the target domain by providing a richer, more contextually diverse set of training data. This leads to the creation of embeddings with rich contextual information.
@@ 3.1.3 Model-Level Adaptation
Model-level adaptation involves developing specialized ViT architectures or modifying existing models to enhance their adaptability to domain shifts. This approach focuses on adapting the internal structure of the model itself to better handle variations between the source and target domains. Techniques such as introducing new layers, modifying attention mechanisms, and designing domain-specific model components are commonly employed in model-level adaptation. These modifications enable the model to learn more robust and transferable features, improving its performance on the target domain.
@@ 3.1.4 Hybrid Approaches
In our categorization of domain adaptation techniques, hybrid approaches combine multiple methods (feature-level, instance-level, and model-level) to leverage the strengths of each. By integrating various techniques, hybrid approaches aim to provide a more robust and comprehensive solution to domain adaptation challenges. These methods effectively address the limitations of individual adaptation strategies by simultaneously aligning feature distributions, selecting relevant instances, and modifying model architectures. Recent studies that fall into our hybrid approaches category include Cross-Domain Vision Transformer (CDTrans), Augmented Transformer, and Multi-View Adaptation. These studies illustrate the effectiveness of hybrid approaches in improving the performance of ViTs across diverse domains.
CDTRANS [87] introduces a hybrid domain adaptation approach via a triple-branch transformer that combines feature-level and model-level adaptations. It incorporates a novel cross-attention module along with self-attention mechanisms within its architecture. The design includes separate branches for source and target data processing and a third for aligning features from both domains. This setup enables simultaneous learning of domain-specific and domain-invariant representations, showing resilience against label noise. Additionally, the paper proposes a two-way center-aware labeling algorithm for the target domain, utilizing a cross-domain similarity matrix to enhance pseudo-label accuracy and mitigate noise impact. This approach effectively merges feature acquisition with alignment, showcasing a sophisticated method for handling domain adaptation challenges. TVT (Transferable Vision Transformer) [88] explores the use of ViT in UDA. It examines ViT’s transferability compared to CNNs and proposes the TVT framework, which includes a Transferability Adaptation Module (TAM) and a Discriminative Clustering Module (DCM). TVT integrates various domain adaptation approaches. It employs model-level adaptations by leveraging ViTs, and optimizing them for UDA. Simultaneously, it involves feature-level adaptation through the TAM, which enhances feature representations for better alignment between source and target domains. Furthermore, instance-level strategies are incorporated via the Discriminative Clustering Module (DCM), focusing on the diversification and discrimination of patch-level features. This multifaceted approach, combining model, feature, and instance-level adaptations, exemplifies a hybrid strategy in domain adaptation.
@@
Category Method Design Highlights Loss Functions Publication
Feature-level Adaptation DOT [90] Domain-Oriented Transformer with dual classifiers utilizes individual tokens for source/target domain adaptation and domain-specific learning, enhanced by source-guided pseudo-label refinement Cross-entropy, Contrastive, Representation Difference ICMR 2022
SUDA [95] Enhancing visual task generalization, leveraging Spectrum Transformer for domain alignment and multi-view learning to optimize mutual information Supervised, Discrepancy, Unsupervised Similarity CVPR 2022
Model-level Adaptation GE-ViT [39] Leveraging ViTs’ inductive biases towards shapes and structures, combined with adversarial and self-supervised learning techniques Cross-Entropy, Entropy, Adversarial CVPR 2022
TFC [102] Integrates CNN and ViTs for both capturing of local and global features, achieving distinct foreground-background separation, further refined by an uncertainty-based accuracy enhancement Cross-Entropy, Adversarial, Uncertainty Penalization IEEE-TCSS 2022
Hybrid Approaches CDTRANS [87] Triple-branch Transformer with shared weights aligns source/target data, using self/cross-attention and center-aware labeling for enhanced pseudo-label accuracy and noise distillation Cross-entropy, Distillation ICLR 2022
TVT [88] Enhances feature learning via adversarial adaptation, using transferability and clustering modules for robust knowledge transfer, and fine-tunes attention for precise domain discrimination Cross-Entropy, Adversarial, Patch-Level Discriminator WACV 2023
SSRT [89] Utilizes adversarial adaptation and enhances model accuracy by refining through perturbed data feedback, incorporating a robust training approach to minimize KL divergence Cross-Entropy, Self-Refinement, Domain Adversarial CVPR 2022
PMTrans [91] Employs a module based on transformer architecture for domain representation, leveraging game theory for patch sampling, and combines source/target data to optimize label accuracy with attention-driven adjustments Cross-Entropy, Semi-Supervised Mixup CVPR 2023
SMTDA [105] Enhances domain adaptation by maximizing nuclear-norm consistency and distilling knowledge via MixUp, refining pseudo labels for broadened generalization across multiple targets Cross-Entropy, Nuclear-Norm, Knowledge Distillation WACV 2023
UniAM [104] Utilizes universal and compressive attention for precise alignment and class separation, achieving categorization without prior label knowledge Cross-Entropy, Adversarial, Source-Target Contrastive ICCV 2023